Dr. Bryan Carpenter

Department of Electronics and Computer Science University of Southampton Highfield Southampton S09 5NH

September 17, 1992

Technical Monographs, Oxford University Computing Laboratory, Programming Research Group, 8-11 Keble Road OXFORD OX1 3QD

Dear Technical Monographs,

I believe you maintain a catalogue (and price-list) for currently available PRG technical reports. Could you please send me a copy of this?

Yours sincerely,